OpenStudy (anonymous):

y varies indirectly as twice x. When x = 1, y = 3. Find y when x = 4. Choose one answer. a. b. c. d.

myininaya (myininaya):

y=k/(2x) 4=k/(2*1) 8=k y=8/(2x) y=4/x y=4/4=1 y=1 when x=4